# Optimal Pure Differentially Private Sparse Histograms in Near-Linear Deterministic Time

### v1

@@ 4.1 Overview of $\mathcal{M}_{\textsc{\scriptsize Hist}}$
For the second requirement, we sample a subset of $n+k-\lvert{I_{1}}\rvert$ random entries from the set $[d]\setminus I_{1}$ (without replacement). This includes not only zero counts, but may also include non-zero ones; i.e., entries in $\operatorname{\textsc{supp}}(\vec{h})$ can have a second chance of being selected. To see how this works, consider one case where $\operatorname{\textsc{supp}}(\vec{h})\cup\{i^{*}\}=\operatorname{\textsc{supp}}(\vec{h}^{\prime})$ for some $i^{*}\in[d]$. In this case, the count $\vec{h}[i^{*}]=0$ and $\vec{h}^{\prime}[i^{*}]=1$. Further, define $I_{1}^{\prime}$ for $\vec{h}^{\prime}$ analogously to $I_{1}$ for $\vec{h}$. Conditioned on $i^{*}\notin I_{1}^{\prime}$ (which happens in most cases), $I_{1}$ and $I_{1}^{\prime}$ have identical distributions. Padding them with sampled entries from the same distribution preserves the distributional identity of the final selected sets.
We only need to handle the case where $i^{*}\in I_{1}^{\prime}$. In this case, $I_{1}$ receives $n+k-\lvert{I_{1}}\rvert$ padding entries, whereas $I_{1}^{\prime}$ receives $n+k-\lvert{I_{1}^{\prime}}\rvert=n+k-\lvert{I_{1}}\rvert-1$ padding entries. The privacy blanket for $I_{1}$ is thus larger by one entry than that for $I_{1}^{\prime}$. Conditioned on this additional entry being $i^{*}$, the remaining parts of the two privacy blankets are identically distributed. Hence, it suffices to show that the probability of $i^{*}$ appearing in the privacy blanket is comparable to its probability of appearing in $I_{1}^{\prime}$, which we will formally establish in Section 4.2.
@@
Up to this point, our discussion naturally raises several questions regarding the proposed padding strategy. For instance, one might wonder what happens if we sample the $n+k-\lvert{I_{1}}\rvert$ padding entries with replacement from the set $[d]\setminus I_{1}$, thereby allowing the final total number of selected entries to be potentially smaller than $n+k$. Another natural question is whether the padding entries could be sampled (with or without replacement) from the larger set $[d]\setminus\operatorname{\textsc{supp}}(\vec{h})$ instead of $[d]\setminus I_{1}$. Do these variations still guarantee pure differential privacy? Do they help avoid re-generating fresh noise for all selected elements in Line 11 of Algorithm 3?
@@
4.2 Proof of Theorem 4.1
In this subsection, we begin the formal proof of Theorem 4.1. We show that Algorithm 3 satisfies the theorem, and analyze its privacy, utility guarantees, and running time separately. All missing proofs of the lemmas in this section can be found in Appendix C.
The noise sampler $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}$ used in Algorithm 3 is described in Theorem 1.2. This mechanism produces noise with the same privacy guarantees and nearly identical tail bounds as discrete Laplace noise, clamped to the range $[0,n]$. For ease of exposition, it is helpful to conceptually treat these samples as clamped discrete Laplace noise. Throughout the proof, we set $k=3\cdot n$,
$\tau\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\;{\big{(}{\Pr\left[1+\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}\left({1}\right)\geq t\right]\leq\varepsilon\gamma/d}\big{)}},$ (2)
and denote $p_{\tau}\doteq\Pr\left[1+\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}\left({1}\right)\geq\tau\right]$.
4.2.1 Privacy Guarantee
@@
$e^{-\max{({\frac{\varepsilon}{2}+p_{\tau},\varepsilon})}}\Pr\left[I=S\right]\leq\Pr\left[I^{\prime}=S\right]\leq e^{\max{\left({\frac{\varepsilon}{2}+p_{\tau}\cdot\frac{d-n-k}{k+1},\varepsilon}\right)}}\cdot\Pr\left[I=S\right],\,\forall\,S\subseteq[d],$ (3)
where $p_{\tau}=\Pr\left[1+{\operatorname{\mathbb{A}px\mathbb{DL}ap}}_{n,\varepsilon,\gamma}\left({1}\right)\geq\tau\right]\leq\varepsilon\gamma/d$.
Conditioned on $I=I^{\prime}$, via the privacy guarantee of $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}$, $\{{\tilde{h}{\left[{i}\right]}:i\in I}\}$ and $\{{\tilde{h}^{\prime}[i]:i\in I^{\prime}}\}$ are $(\varepsilon,0)$-indistinguishable. By basic composition theorem of DP (A.1), the Algorithm 3 is $\max{\left({\frac{3\varepsilon}{2}+p_{\tau}\cdot\frac{d-n-k}{k+1},2\varepsilon}\right)}=2$-DP. It remains to prove Equation 3.
Proving Equation 3. Without lose of generality, we assume that $\vec{h}$ and $\vec{h}^{\prime}$ differ in two entries indexed by $i^{*},j^{*}\in[d]$. Further, we assume that
$\begin{array}[]{ccc}\vec{h}{\left[{i^{*}}\right]}+1=\vec{h}^{\prime}[i^{*}],&\vec{h}{\left[{j^{*}}\right]}=\vec{h}^{\prime}[j^{*}]+1,&\vec{h}{\left[{i}\right]}=\vec{h}^{\prime}[i],\,\forall i\in[d]\setminus\{{i^{*},j^{*}}\}.\end{array}$ (5)
@@
Case 1. $\operatorname{\textsc{supp}}{({\vec{h}})}=\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}$. Then clearly the distributions of $I_{1}$ and $I_{1}^{\prime}$ are $(\varepsilon,0)$-indistinguishable, due to the privacy guarantee of $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}$. Hence so are the ones of $I$ and $I^{\prime}$ and Equation 3 holds.
Case 2. $\operatorname{\textsc{supp}}{({\vec{h}})}\cup\{{i^{*}}\}=\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}$. In this case, we have
First, observe that
$\Pr\left[I=S\right]$ $=\sum_{J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}}\Pr\left[I=S\mid I_{1}=J\right]\cdot\Pr\left[I_{1}=J\right],$
$\Pr\left[I^{\prime}=S\right]$ $=\sum_{J^{\prime}\subseteq\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}}\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J^{\prime}\right]\cdot\Pr\left[I_{1}^{\prime}=J^{\prime}\right].$
Therefore
$\Pr\left[I^{\prime}=S\right]$ $=\sum_{J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}}\Big{(}\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\right]\Pr\left[I_{1}^{\prime}=J\right]+\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]\Pr\left[I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]\Big{)}.$
We trivially see that $\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\right]=\Pr\left[I=S\mid I_{1}=J\right]$ by on Algorithm 3. To show that $\Pr\left[I=S\right]$ and $\Pr\left[I^{\prime}=S\right]$ are close, we need to investigate the other quantities involved in their sums, as demonstrated by the following lemmas.
@@
Observe that, by the privacy guarantee of the noise sampler used in Algorithm 3 (Theorem 5.2), it holds that $r_{J}\in(e^{-\varepsilon/2},e^{\varepsilon/2})$.
@@
Let $\kappa\doteq\frac{d-n+1}{k+1}$. For each $J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}$, it holds that
$\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]\leq\kappa\cdot\Pr\left[I=S\mid I_{1}=J\right].$ (6)
Based on Lemma 4.2 and that $\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\right]=\Pr\left[I=S\mid I_{1}=J\right]$, we have
$\Pr\left[I^{\prime}=S\right]$ $={({1-p_{\tau}})}\cdot\sum_{J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}}r_{J}\cdot{\Pr\left[I=S\mid I_{1}=J\right]\cdot\Pr\left[I_{1}=J\right]}$
$+\sum_{J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}}r_{J}\cdot{\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]\cdot p_{\tau}\cdot\Pr\left[I_{1}=J\right]}.$
We are now ready to prove Equation 3.
Proving LHS of Equation 3: since $r_{J}\geq e^{-\varepsilon/2}$ for all $J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}$, it is easy to see that
$\Pr\left[I^{\prime}=S\right]$ $\geq{({1-p_{\tau}})}\cdot e^{-\varepsilon/2}\cdot\sum_{J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}}\Pr\left[I=S\mid I_{1}=J\right]\cdot\Pr\left[I_{1}=J\right]$ (7)
$\geq{({1-p_{\tau}})}\cdot e^{-\varepsilon/2}\cdot\Pr\left[I=S\right]\geq e^{-2\cdot p_{\tau}-\varepsilon/2}\cdot\Pr\left[I=S\right].$ (8)
where $1-p_{\tau}\geq e^{-2\cdot p_{\tau}}$ holds when $p_{\tau}\in{({0,0.5})}$.
Proving RHS of Equation 3: since $r_{J}\leq e^{\varepsilon/2}$ for all $J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}$, and based on Lemma 4.3
$\Pr\left[I^{\prime}=S\right]$ $\leq e^{\varepsilon/2}\cdot{\left({{({1-p_{\tau}})}\cdot\Pr\left[I=S\right]+p_{\tau}\cdot\kappa\cdot\sum_{J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}}\Pr\left[I=S\mid I_{1}=J\right]\cdot\Pr\left[I_{1}=J\right]}\right)}$
$=e^{\varepsilon/2}\cdot{\left({{({1-p_{\tau}})}+p_{\tau}\cdot\kappa}\right)}\cdot\Pr\left[I=S\right]$
$=e^{\varepsilon/2}\cdot{\left({1+p_{\tau}\cdot{({\kappa-1})}}\right)}\cdot\Pr\left[I=S\right]$
$\leq e^{\varepsilon/2+p_{\tau}\cdot{({\kappa-1})}}\cdot\Pr\left[I=S\right].$
Observing $\kappa-1=\frac{d-n+1-{\left({k+1}\right)}}{k+1}=\frac{d-n-k}{k+1}$ proves the RHS of Equation 3.
Case 3. $\operatorname{\textsc{supp}}{({\vec{h}})}=\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}\cup\{{j^{*}}\}$. The discussion follows from the symmetry of Case 2.
Case 4. $\operatorname{\textsc{supp}}{({\vec{h}})}\cup\{{i^{*}}\}=\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}\cup\{{j^{*}}\}$. First, observe that the subsets in $\operatorname{\textsc{supp}}{({\vec{h}})}$ can be partitioned into those containing $j^{*}$ and those that do not. For each $J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}$ with $j^{*}\notin J$, we have $J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}\cap\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}$. Hence,
$\Pr\left[I=S\right]$ $=\sum_{J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}\cap\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}}{\Pr\left[I=S\mid I_{1}=J\right]\cdot\Pr\left[I_{1}=J\right]}$ (9)
Similarly,
$\Pr\left[I^{\prime}=S\right]$ $=\sum_{J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}\cap\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}}{\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\right]\cdot\Pr\left[I_{1}^{\prime}=J\right]}$ (10)
For each $J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}\cap\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}$, define
$q_{J}$ $\doteq\Pr\left[I=S\mid I_{1}=J\right]\cdot\Pr\left[I_{1}=J\right]+\Pr\left[I=S\mid I_{1}=J\cup\{{j^{*}}\}\right]\cdot\Pr\left[I_{1}=J\cup\{{j^{*}}\}\right],$
$q_{J}^{\prime}$ $\doteq\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\right]\cdot\Pr\left[I_{1}^{\prime}=J\right]+\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]\cdot\Pr\left[I_{1}^{\prime}=J\cup\{{i^{*}}\}\right].$
We want to compare $q_{J}$ and $q_{J}^{\prime}$. Our goal is to prove that
$e^{-\frac{d-n+1}{k+1}\cdot p_{\tau}}\cdot q_{J}\leq q_{J}^{\prime}\leq e^{\frac{d-n+1}{k+1}\cdot p_{\tau}}\cdot q_{J},$ (11)
which immediately gives
$e^{-\frac{d-n+1}{k+1}\cdot p_{\tau}}\cdot\Pr\left[I=S\right]\leq\Pr\left[I^{\prime}=S\right]\leq e^{\frac{d-n+1}{k+1}\cdot p_{\tau}}\cdot\Pr\left[I=S\right].$
Proving Equation 11. Clearly we have $\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\right]=\Pr\left[I=S\mid I_{1}=J\right]$, which follows directly from the way Algorithm 3 works. To proceed, we need the following lemmas.
@@
Assume that $\operatorname{\textsc{supp}}{({\vec{h}})}\cup\{{i^{*}}\}=\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}\cup\{{j^{*}}\}$. Then for any $J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}\cap\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}$, it holds that
$\Pr\left[I_{1}^{\prime}=J\right]=\Pr\left[I_{1}=J\right],\quad\Pr\left[I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]=\Pr\left[I_{1}=J\cup\{{j^{*}}\}\right].$
There are four cases to be discussed to prove Equation 11.
Case (i): $J\nsubseteq S$, then $q_{J}=q_{J}^{\prime}=0$ since
$\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]=\Pr\left[I=S\mid I_{1}=J\cup\{{j^{*}}\}\right]=0.$
Case (ii): $J\subseteq S$ and $i^{*}\in S$, $j^{*}\in S$, then $q_{J}=q_{J}^{\prime}$ since due to symmetry we have
$\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]$ $=\Pr\left[I_{2}^{\prime}=S\setminus{\left({J\cup\{{i^{*}}\}}\right)}\right]$
$=\Pr\left[I_{2}=S\setminus{\left({J\cup\{{j^{*}}\}}\right)}\right]=\Pr\left[I=S\mid I_{1}=J\cup\{{j^{*}}\}\right].$
Case (iii): $J\subseteq S$ and $i^{*}\notin S$, $j^{*}\in S$. We have the following lemma.
Lemma 4.5.
Denote $\kappa^{\prime}=\frac{d-\lvert{J}\rvert}{n+k-\lvert{J}\rvert}$. Then $\frac{\Pr\left[I=S\mid I_{1}=J\cup\{{j^{*}}\}\right]missing}{\Pr\left[I=S\mid I_{1}=J\right]missing}=\kappa^{\prime}$.
Based on the Lemma 4.5 and that $\Pr\left[I_{1}=J\cup\{{j^{*}}\}\right]=p_{\tau}\cdot\Pr\left[I_{1}=J\right]$,
$q_{J}$ $=\Pr\left[I=S\mid I_{1}=J\right]\cdot\Pr\left[I_{1}=J\right]+\Pr\left[I=S\mid I_{1}=J\right]\cdot\kappa^{\prime}\cdot p_{\tau}\cdot\Pr\left[I_{1}=J\right].$
$={\left({1+\kappa^{\prime}\cdot p_{\tau}}\right)}\cdot\Pr\left[I=S\mid I_{1}=J\right]\cdot\Pr\left[I_{1}=J\right].$
Since $\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]$ in this case,
$q_{J}^{\prime}$ $=\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\right]\cdot\Pr\left[I_{1}^{\prime}=J\right]=\Pr\left[I=S\mid I_{1}=J\right]\cdot\Pr\left[I_{1}=J\right].$
Therefore,
$q_{J}^{\prime}$ $\leq q_{J}\leq\exp{\left({\kappa^{\prime}\cdot p_{\tau}}\right)}\cdot q_{J}^{\prime}\leq e^{\frac{d-n+1}{k+1}\cdot p_{\tau}}\cdot q_{J}^{\prime}.$
Case (iv): $J\subseteq S$ and $i^{*}\in S$, $j^{*}\notin S$, this is a symmetric to Case Two.
4.2.2 Utility Guarantee

### v2

@@ 4.1. Overview of $\mathcal{M}_{\textsc{\scriptsize Hist}}$
For the second requirement, we sample a subset of $n+k-\lvert{I_{1}}\rvert$ random entries from the set $d]\setminus I_{1}$ (without replacement). This includes not only zero counts, but may also include non-zero ones; i.e., entries in $\operatorname{\textsc{supp}}(\vec{h})$ can have a second chance of being selected. To see how this works, consider one case where $\operatorname{\textsc{supp}}(\vec{h})\cup\{i^{*}\}=\operatorname{\textsc{supp}}(\vec{h}^{\prime})$ for some $i^{*}\in[d]$. In this case, the count $\vec{h}[i^{*}]=0$ and $\vec{h}^{\prime}[i^{*}]=1$. Further, define $I_{1}^{\prime}$ for $\vec{h}^{\prime}$ analogously to $I_{1}$ for $\vec{h}$. Conditioned on $i^{*}\notin I_{1}^{\prime}$ (which happens in most cases), $I_{1}$ and $I_{1}^{\prime}$ have identical distributions. Padding them with sampled entries from the same distribution preserves the distributional identity of the final selected sets.
We only need to handle the case where $i^{*}\in I_{1}^{\prime}$. In this case, $I_{1}$ receives $n+k-\lvert{I_{1}}\rvert$ padding entries, whereas $I_{1}^{\prime}$ receives $n+k-\lvert{I_{1}^{\prime}}\rvert=n+k-\lvert{I_{1}}\rvert-1$ padding entries. The privacy blanket for $I_{1}$ is thus larger by one entry than that for $I_{1}^{\prime}$. Conditioned on this additional entry being $i^{*}$, the remaining parts of the two privacy blankets are identically distributed. Hence, it suffices to show that the probability of $i^{*}$ appearing in the privacy blanket is comparable to its probability of appearing in $I_{1}^{\prime}$, which we will formally establish in Section˜4.2.
@@
Up to this point, our discussion naturally raises several questions regarding the proposed padding strategy. For instance, one might wonder what happens if we sample the $n+k-\lvert{I_{1}}\rvert$ padding entries with replacement from the set $d]\setminus I_{1}$, thereby allowing the final total number of selected entries to be potentially smaller than $n+k$. Another natural question is whether the padding entries could be sampled (with or without replacement) from the larger set $d]\setminus\operatorname{\textsc{supp}}(\vec{h})$ instead of $d]\setminus I_{1}$. Do these variations still guarantee pure differential privacy? Do they help avoid re-generating fresh noise for all selected elements in Line 11 of Algorithm˜2?
@@
4.2. Proof of Theorem 4.1
In this subsection, we begin the formal proof of Theorem˜4.1. We show that Algorithm˜2 satisfies the theorem, and analyze its privacy, utility guarantees, and running time separately. All missing proofs of the lemmas in this section can be found in Appendix˜C.
Throughout the proof, we set
(2) $\tau\doteq\operatorname*{arg\,min}_{t\in\mathbb{N}}\left\{{\Pr\left[1+\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}\left({1}\right)\geq t\right]\leq 1/d}\right\},$
Based on Theorem 1.4, it holds that $\tau\in\Theta{({1/\varepsilon\cdot\ln d})}$. Further, denote
(3) $p_{\tau}\doteq\Pr\left[1+\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}\left({1}\right)\geq\tau\right].$
The definition of $\tau$ implies that $p_{\tau}\leq 1/d$.
4.2.1. Privacy Guarantee
@@
(4) $e^{-\max{({\frac{\varepsilon}{2}+2\cdot p_{\tau},\varepsilon})}}\Pr\left[I=S\right]\leq\Pr\left[I^{\prime}=S\right]\leq e^{\max{\left({\frac{\varepsilon}{2}+p_{\tau}\cdot\frac{d-k}{k+1},\varepsilon}\right)}}\cdot\Pr\left[I=S\right],\,\forall\,S\subseteq[d].$
Conditioned on $I=I^{\prime}$, by the privacy guarantee of $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}$ (Theorem 1.4), the values $\{{\tilde{h}{\left[{i}\right]}:i\in I}\}$ and $\{{\tilde{h}^{\prime}[i]:i\in I^{\prime}}\}$ are $(\varepsilon,0)$-indistinguishable. By basic composition theorem of DP (˜A.1) and that $p_{\tau}\leq 1/d$, the Algorithm˜2 is
$\max{\Big({\frac{3\varepsilon}{2}+p_{\tau}\cdot\frac{d-k}{k+1},2\varepsilon}\Big)}\leq\max{\Big({\frac{3\varepsilon}{2}+\frac{1}{k+1},2\varepsilon}\Big)}\text{-DP}$
It remains to prove Equation˜4.
Proving Equation˜4.
Without lose of generality, we assume that $\vec{h}$ and $\vec{h}^{\prime}$ differ in two entries indexed by $i^{*},j^{*}\in[d]$. Further, we assume that
$\begin{array}[]{ccc}\vec{h}{\left[{i^{*}}\right]}+1=\vec{h}^{\prime}[i^{*}],&\vec{h}{\left[{j^{*}}\right]}=\vec{h}^{\prime}[j^{*}]+1,&\vec{h}{\left[{i}\right]}=\vec{h}^{\prime}[i],\,\forall i\in[d]\setminus\{{i^{*},j^{*}}\}.\end{array}$
@@
Case 1. $\operatorname{\textsc{supp}}{({\vec{h}})}=\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}$
Then clearly the distributions of $I_{1}$ and $I_{1}^{\prime}$ are $(\varepsilon,0)$-indistinguishable, due to the privacy guarantee of $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}$. Hence so are the ones of $I$ and $I^{\prime}$ and Equation˜4 holds.
Case 2. $\operatorname{\textsc{supp}}{({\vec{h}})}\cup\{{i^{*}}\}=\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}$.
In this case, we have $\vec{h}^{\prime}[i^{*}]=1,\,\vec{h}{\left[{i^{*}}\right]}=0,\,\vec{h}{\left[{j^{*}}\right]}>\vec{h}^{\prime}[j^{*}]>0.$ For convenience, denote $\mu$ as the distribution of $I$, i.e., $\mu(S)=\Pr\left[I=S\right]$. Similarly, let $\mu^{\prime}$ be the distribution of $I^{\prime}$, with $\mu^{\prime}(S)=\Pr\left[I^{\prime}=S\right]$. Let $\mathcal{E}$ be the event that $i^{*}\in I_{1}^{\prime}$, and let $\mu^{\prime}_{\mathcal{E}}$ and $\mu^{\prime}_{\bar{\mathcal{E}}}$ denote the distributions of $I^{\prime}$ conditioned on $\mathcal{E}$ and $\bar{\mathcal{E}}$, respectively. By the definition of $p_{\tau}$,
$\mu^{\prime}=p_{\tau}\cdot\mu^{\prime}_{\mathcal{E}}+{({1-p_{\tau}})}\cdot\mu^{\prime}_{\bar{\mathcal{E}}}.$
First, to prove the LHS of Equation˜4, we invoke Lemma 4.2.
@@
$\mu$ and $\mu^{\prime}_{\bar{\mathcal{E}}}$ are $({\varepsilon/2,0})$}-indistinguishable.
Intuitively, the lemma holds because, conditioned on the event $\bar{\mathcal{E}}$ (i.e., $i^{*}\notin I_{1}^{\prime}$), the distributions of $I_{1}^{\prime}$ and $I_{1}$ become $({\varepsilon/2,0})$}-indistinguishable: the only remaining difference between $\vec{h}$ and $\vec{h}^{\prime}$ is the single-coordinate change $\vec{h}[j^{*}]=\vec{h}^{\prime}[j^{*}]+1$, and the privacy guarantee of $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}$ ensures this. Moreover, conditioned on $I_{1}^{\prime}=I_{1}$, the distributions of $I_{2}^{\prime}$ and $I_{2}$ coincide. Consequently, $\mu^{\prime}_{\bar{\mathcal{E}}}$ and $\mu$ are $({\varepsilon/2,0})$}-indistinguishable.
By Lemma˜4.2,
$\mu^{\prime}$ $\geq{({1-p_{\tau}})}\cdot\mu^{\prime}_{\bar{\mathcal{E}}}\geq{({1-p_{\tau}})}\cdot e^{-\varepsilon/2}\cdot\mu\geq e^{-2p_{\tau}-\varepsilon/2}\cdot\mu,$
where we use that $1-p_{\tau}\geq e^{-2p_{\tau}}$ for $p_{\tau}\in(0,0.5)$.
Second, to prove the RHS of Equation˜4, we use Lemma 4.3.
@@
$\mu\geq\frac{k}{d}\cdot e^{-\varepsilon/2}\cdot\mu^{\prime}_{\mathcal{E}}$.
Informally, the lemma holds because, the element $i^{*}$ can be added to $I$ only during the padding phase (Algorithm 2, Lines 89), which occurs with probability at least $k/d$. Conditioned on this event and on $\mathcal{E}$, the distributions of $I_{1}\cup\{{i^{*}}\}$ and $I_{1}^{\prime}$ are $({\varepsilon/2,0})$}-indistinguishable: the only remaining difference between $\vec{h}$ and $\vec{h}^{\prime}$ is the single-coordinate change $\vec{h}[j^{*}]=\vec{h}^{\prime}[j^{*}]+1$, and the privacy guarantee of $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}$ ensures this. Moreover, when $I_{1}\cup\{{i^{*}}\}=I_{1}^{\prime}$ and $i^{*}\in I_{2}$, the remaining padding elements in $I_{2}\setminus\{{i^{*}}\}$ and $I_{2}^{\prime}$ have identical distributions.
Based on Lemmas˜4.2 and 4.3,
$\mu^{\prime}$ $=p_{\tau}\cdot\mu^{\prime}_{\mathcal{E}}+{({1-p_{\tau}})}\cdot\mu^{\prime}_{\bar{\mathcal{E}}}\leq p_{\tau}\cdot e^{\varepsilon/2}\cdot d/k\cdot\mu+{({1-p_{\tau}})}\cdot e^{\varepsilon/2}\cdot\mu$
$=e^{\varepsilon/2}\cdot{\left({{({1-p_{\tau}})}+p_{\tau}\cdot d/k}\right)}\cdot\mu=e^{\varepsilon/2}\cdot{\left({1+p_{\tau}\cdot{({d/k-1})}}\right)}\cdot\mu\leq e^{\varepsilon/2+p_{\tau}\cdot{({d/k-1})}}\cdot\mu,$
where we use $1+p_{\tau}\cdot{({d/k-1})}\leq\exp{({p_{\tau}\cdot{({d/k-1})}})}$.
Case 3. $\operatorname{\textsc{supp}}{({\vec{h}})}=\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}\cup\{{j^{*}}\}$.
The discussion follows from the symmetry of Case 2.
Case 4. $\operatorname{\textsc{supp}}{({\vec{h}})}\cup\{{i^{*}}\}=\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}\cup\{{j^{*}}\}$.
Let $\mu$, $\mu^{\prime}$, $\mu^{\prime}_{\mathcal{E}}$ and $\mu^{\prime}_{\bar{\mathcal{E}}}$ be defined as before. Further, let $\mathcal{F}$ be the event that $j^{*}\in I_{1}$, and let $\mu_{\mathcal{F}}$ and $\mu_{\bar{\mathcal{F}}}$ denote the distributions of $I$ conditioned on $\mathcal{F}$ and $\bar{\mathcal{F}}$, respectively. By the definition of $p_{\tau}$,
$\mu=p_{\tau}\cdot\mu_{\mathcal{F}}+{({1-p_{\tau}})}\cdot\mu_{\bar{\mathcal{F}}},\quad\mu^{\prime}=p_{\tau}\cdot\mu^{\prime}_{\mathcal{E}}+{({1-p_{\tau}})}\cdot\mu^{\prime}_{\bar{\mathcal{E}}}.$
To prove Equation˜4, we need Lemma 4.4.
@@
$\mu_{\bar{\mathcal{F}}}=\mu^{\prime}_{\bar{\mathcal{E}}}$, $\mu_{\bar{\mathcal{F}}}\geq\frac{k}{d}\cdot\mu^{\prime}_{\mathcal{E}}$, $\mu^{\prime}_{\bar{\mathcal{E}}}\geq\frac{k}{d}\cdot\mu_{\mathcal{F}}$.
We present an intuitive explanation of the lemma. First, conditioned on $\bar{\mathcal{F}}$ and $\bar{\mathcal{E}}$, the pair $I_{1}$ and $I_{1}^{\prime}$, as well as the pair $I_{2}$ and $I_{2}^{\prime}$, have identical distributions. Therefore, $\mu_{\bar{\mathcal{F}}}=\mu^{\prime}_{\bar{\mathcal{E}}}$ holds. Second, the element $i^{*}$ can be added to $I$ only during the padding phase (Algorithm 2, Lines 89), which occurs with probability at least $k/d$. Conditioned on this event and on $\bar{\mathcal{F}}$ and $\mathcal{E}$, the pair $I_{1}$ and $I_{1}^{\prime}$, as well as the pair $I_{2}$ and $I_{2}^{\prime}$, have identical distributions. Therefore, $\mu_{\bar{\mathcal{F}}}\geq\frac{k}{d}\cdot\mu^{\prime}_{\mathcal{E}}$ holds. The last inequality follows from symmetry of the second one.
Based on Lemma˜4.4
$\mu$ $=p_{\tau}\cdot\mu_{\mathcal{F}}+{({1-p_{\tau}})}\cdot\mu_{\bar{\mathcal{F}}}\leq p_{\tau}\cdot\tfrac{d}{k}\cdot\mu^{\prime}_{\bar{\mathcal{E}}}+{({1-p_{\tau}})}\cdot\mu^{\prime}_{\bar{\mathcal{E}}}\leq{\left({1+p_{\tau}\cdot{\Big({\frac{d}{k}-1}\Big)}}\right)}\cdot\mu^{\prime}\leq e^{\,p_{\tau}\cdot\frac{d-k}{k}}\mu^{\prime}.$
4.2.2. Utility Guarantee